OpenStudy (anonymous):

The band booster and the athletic booster clubs sell tickets for a fund-raiser. The band booster sells eight more than twice the number of tickets the athletic booster sells. The band booster club sells 280 tickets. Find the number of tickets the athletic booster club sells. a. 144 tickets c. 136 tickets b. 132 tickets d. 148 tickets

OpenStudy (anonymous):

280=8+2x 280-8=272 272/2=136 so the answer is 136